1) how many gallons each of 25% alcohol and 35% alcohol should be mixed to get 20 gallons of 32% alcohol?
Answer by ewatrrr(24785)

 
Hi,
mixed to get 20 gallons
Let x and (20-x) represent the amount of the 35% and 25% solutions respectively
Question states***
mixing to get 20 gal of 32% alcohol
.35x + .25(20-x) = .32*20G
Solving for x
.10x = .07*20G =
x = 14G, the amount of the 35% Solution. 6G of the 25% solution
CHECKING our Answer***
.35*14 + .25*6 = 4.9 + 1.5 = .32*20 = .64